OpenStudy (anonymous):

solve by substitution method 9x+7y=-24 x=19-8y

OpenStudy (anonymous):

putting x=19-8y in first equation 9(19-8y)+7y=-24 171-72y+7y=-24 -65y=-195 y=3 putting y=3 in second eqn x=19-(8*3) x=19-24=-5 y=3, x=-5

OpenStudy (anonymous):

THAT'S right